OpenStudy (anonymous):

Simplifying Radicals a^1/2/a^2

OpenStudy (haseeb96):

@taramgrant0543664 can you help this also ? LOL

OpenStudy (unklerhaukus):

\[\frac{a^n}{a^m}=a^{n-m}\]

OpenStudy (xapproachesinfinity):

You got to it first #unkle lol

OpenStudy (xapproachesinfinity):

, about to post same thing

OpenStudy (unklerhaukus):

\[a^{-p}=\frac1{a^p}\]

OpenStudy (unklerhaukus):

what is your final answer @Heartz

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\large \cfrac{a^n}{a^m}\implies a^{n-m}\qquad thus\qquad \cfrac{a^{\frac{1}{2}}}{a^2}\implies a^{\frac{1}{2}-2}\implies a^{\square ?}\)

OpenStudy (jacob902):

a^3/2

OpenStudy (jdoe0001):

well... notice,the "2" is negative though

OpenStudy (jdoe0001):

\(\large { \cfrac{a^n}{a^m}\implies a^{n-m}\qquad thus\qquad \cfrac{a^{\frac{1}{2}}}{a^2}\implies a^{\frac{1}{2}-2}\qquad \begin{cases} 2=\cfrac{4}{2} \end{cases}\qquad thus \\ \quad \\ a^{\frac{1}{2}-2}\implies a^{\frac{1}{2}-\frac{4}{2}}\implies a^{\square ?} }\)